# Stronger ViTs With Octic Equivariance

### v2

@@ 6 Conclusion
We introduced octic-equivariant ViT layers that, when incorporated, maintain or improve representational power while significantly reducing computational complexity. We validated our proposed architectures by their effectiveness in both supervised and self-supervised learning. In particular, we achieved approximately 40% reduction in FLOPs for ViT-H while consistently improving performance, positioning octic ViTs as a strong addition to the catalog of vision architectures.
@@ 31 SQRT2_OVER_4 * (bmd + emh)
Appendix B Visualizations
B.1 Octic Layers
We visualize how the features are transformed in some of the layers in Figure 5. All layers take in $\frac{D}{8}\rho_{\text{iso}}$ features and output features of the same type. For the linear layer, we use five different weight matrices, one for each irrep, and where weights are shared among the two dimensions of the E features. In the Attention layer, we project the features to query, key and values and flatten them along the final dimension. We can then apply standard multi-head self-attention (MHA) using FlashAttention [15]. Thereafter, we divide the features into the same format as the input to the layer. Non-linearities such as GELU can be applied after transforming the features from $\rho_{\text{iso}}$ to $\rho_{\text{reg}}$ (through an inverse Fourier transform) and then back to $\rho_{\text{iso}}$.
Figure 5: Octic Layers. Illustration of how the features are combined in the Linear (left), Attention (middle) and GELU (right) layers.
B.2 Learned Octic Embeddings
The PatchEmbed layer contains filters mapping the input from 3 channel dimensions to $D$ embedding dimensions. To illustrate the learned filters we take inspiration from Dosovitskiy et al. [21] and visualize the first 16 principal components. In contrast to regular ViTs, we have six different learned filters corresponding to the five irreps. Four for the one dimensional and two for the E irrep (due to its multiplicity). The results are illustrated in Figure 6. Interestingly, the learned filters look qualitatively different between the two learning methods. This is similar to how the learned filters of baseline DINOv2 and DeiT III look qualitatively different. DINOv2 training appears to produce more high frequency patterns while DeiT III gives clearer patterns. For the invariant irrep (A1), it appears as though DeiT III training produces a spherical pattern.
@@ B.2 Learned Octic Embeddings
Appendix C Invariantization
There are multiple options to produce $\mathrm{D}_{8}$ invariant features, i.e. mapping tokens of type $\rho_{\text{chan}}=\frac{D}{8}\rho_{\text{iso}}$ to features of type $D\rho_{\text{A1}}$ (here denoted in short as invariantization). We let $\psi$ be a function mapping from features of type $\frac{D}{8}\rho_{\text{iso}}$ to features of type $\frac{KD}{8}\rho_{\text{A1}}$ for some $K$ which can be larger or smaller than $8$. These $KD/8$ dimensions are then mapped through a small MLP to $D$ dimensions again.
@@ Appendix C Invariantization
The triple correlation method [43, 29] extracts a complete set of third order homogeneous polynomial invariants from a signal over $\mathrm{D}_{8}$. We computed a basis for all third order invariant homogeneous polynomials using Macaulay2 [26, 24] and use the basis elements as invariant. Here, $K=15$.
@@
We follow Cahill et al. [9] and implement a version of their max filtering invariant. For this, we have a set of $2D$ learnable $D$-dimensional tokens $\mathbf{y}\in\mathbb{R}^{2D\times D}$, and the $2D$ invariants are given by
$\psi(x)=\oplus_{k=1}^{2D}\max_{g\in\mathrm{D}_{8}}\langle\mathbf{y}_{k},\rho_{\text{chan}}(g)x\rangle.$ (17)
@@
Similar to the max filtering approach, we implement a canonisation where we have a single learnable $D$-dimensional reference token $y$ and compute the $D$-dimensional invariant as
@@
[…]
@@
Appendix D Experimental Setting
D.1 DeiT III
We train for 400 epochs on ImageNet-1K with an effective batch size of 2048 following Touvron et al. [46] and Bökman et al. [8]. We compare to the figures reported in the respective papers and thus only train the octic ViTs. The training recipe includes heavy data augmentation (e.g. cutmix, mixup and color jitter) and uses the deprecated NVIDIA Apex library. Training is done in mixed precision with the lamb optimizer. The most important hyperparameters are summarized in Table 5(a). For more details, we refer to [46, 8] and the official repo used for reproduction https://github.com/facebookresearch/deit.
@@ D.1 DeiT III
Init layer scale $1.10^{-5}$ $1.10^{-5}$
@@
D.2 DINOv2
We closely follow the implementation in the original paper, only modifying to train in BF16 instead of FP16 for greater stability, and follow the same evaluation protocol for classification. Lacking an official reproduction of the segmentation protocol in DINOv2, we opt for the evaluation protocol created by Darcet et al. [16] for semantic segmentation on ADE20K and VOC2012. In contrast to the original DINOv2 paper, we decide to limit our study to ViT-L and ViT-H, the latter of which was not included in the original paper. We opt for the larger patch size of $P=16$ for all our DINOv2 models to save computational resources. Note, this is the reason why we report fewer FLOPs for our DINOv2 ViT-H models than their DeiT III counterpart (which use $P=14$).
@@ D.2 DINOv2
D.3 DinoBloom
@@
For the details of the experiment, we closely follow the training and evaluation protocol of Koch et al. [32]. In particular, we finetune our DINOv2 checkpoints for 4K iterations (taking approx. 1 hour on 8 A100-40GB) and evaluate on a hold-out split of the Bone Marrow Cytomorphology (BMC) [35] dataset. However, we limit our finetuning datasets to the datasets presented in Table 6. Note, we follow the same datasplit of BMC as in DinoBloom and thus also refrain from training on that part.
@@
[…]
D.4 General Settings
@@
We utilize the PyTorch [38] and the timm [54] libraries for our experiments. We run the same versioning as our benchmarks. For all other experiments, we use Python 3.11.9 and PyTorch 2.6.0 with CUDA 11.8.
@@
The model sizes referred to in the paper adhere to the standard terminology used by Wightman [54], Dosovitskiy et al. [21]. If we denote the shape by a tuple of (depth, width, attention heads), ViT-L has shape (24, 1024, 16) and ViT-H has shape (32, 1280, 16). Both use MLP dimension four times the size of the embedding dimension (commonly referred to as MLP ratio).
@@
Throughput and peak memory are measured on a single A100-80GB GPU with batch size fixed to 64 using torch.compile, FlashAttention [15], and mixed precision. The throughput only measures forward passes with no gradients. Moreover, we utilize 10 warm-up iterations and then average over 100 runs [8]. Peak memory is measured with PyTorch’s device memory allocation monitor.
@@
We count the number of FLOPs using fvcore.nn.FlopCountAnalysis https://github.com/facebookresearch/fvcore. FLOPs are normalized with respect to the batch size (i.e. we measure FLOPs/image). We acknowledge that the term FLOPs often leads to confusion. We adopt the terminology of prior work [46, 8] and the fvcore library for FLOPs, though, strictly speaking, this refers to MACs (as a factor of two is omitted).
D.5 Compute Resources
Table 8 provides information on the computing resources required for the main experiments. The largest single experiment (i.e. not accounting ablation studies) used 32×32$\times$32 ×A100-40GB for 58 hours. Failed or discarded results are not included. In total, we used around 20k A100-40GB hours for our main results.
@@ D.5 Compute Resources
DeiT III $\mathcal{I}_{8}$(ViT-H/14) 32×32$\times$32 ×A100-40GB 58h 1856
DeiT III $\mathrm{D}_{8}$(ViT-H/14) 32×32$\times$32 ×A100-40GB 57h 1824
DeiT III $\mathcal{H}_{8}$(ViT-H/14) 32×32$\times$32 ×A100-40GB 55h 1760
DeiT III $\mathcal{H}_{8}$(ViT-L/16) 16×16$\times$16 ×A100-40GB 39h 624
DeiT III $\mathcal{I}_{8}$(ViT-L/16) 16×16$\times$16 ×A100-40GB 37h 592
DeiT III $\mathrm{D}_{8}$(ViT-L/16) 16×16$\times$16 ×A100-40GB 36h 576
Ablation: Invarisation $\mathcal{H}_{8}$(ViT-L/16) 16×16$\times$16 ×A100-40GB 240h 3840
Ablation: Hybridisation $\mathcal{I}_{8}$(ViT-L/16) 16×16$\times$16 ×A100-40GB 122h 1952
Ablation: Hybridisation $\mathcal{H}_{8}$(ViT-L/16) 16×16$\times$16 ×A100-40GB 120h 1920
DINOv2 $\mathcal{H}_{8}$(ViT-H/16) 32×32$\times$32 ×A100-40GB 33h 1056
DINOv2 $\mathcal{I}_{8}$(ViT-H/16) 32×32$\times$32 ×A100-40GB 30h 960
DINOv2 $\mathrm{D}_{8}$ (ViT-L/16) 32×32$\times$32 ×A100-40GB 23h 736
DINOv2 ViT-H/16 32×32$\times$32 ×A100-40GB 21h 672
DINOv2 $\mathcal{I}_{8}$(ViT-L/16) 32×32$\times$32 ×A100-40GB 19h 608
DINOv2 $\mathcal{H}_{8}$(ViT-L/16) 32×32$\times$32 ×A100-40GB 19h 608
DINOv2 ViT-L/16 16×16$\times$16 ×A100-40GB 24h 384
DinoBloom $\mathcal{I}_{8}$(ViT-L/16) 8×8$\times$8 ×A100-40GB 46min 6
DinoBloom ViT-L/16 8×8$\times$8 ×A100-40GB 37min 5
@@ Appendix E Licenses
All our models are trained on ImageNet-1K [18] which has a custom non-commercial license (see https://www.image-net.org/download.php). We further evaluate on ADE20K, VOC2012, and BMC. For the licenses see https://ade20k.csail.mit.edu/terms/index.html, http://host.robots.ox.ac.uk/pascal/VOC/voc2012/, and https://www.cancerimagingarchive.net/data-usage-policies-and-restrictions/, respectively.
@@
The efficient octic layers are our original work and will be licensed under Apache License 2.0 following the code release. The training pipelines are from DeiT III [46] and DINOv2 [37] which are both also under the Apache License 2.0. For details about the Apache License 2.0 we refer to https://www.apache.org/licenses/LICENSE-2.0.

### v4

@@ 6 Conclusion
We introduced octic-equivariant ViT layers that, when incorporated, maintain accuracy while significantly reducing computational complexity. We validated our proposed architectures by their effectiveness in both supervised and self-supervised learning, and conducted ablation studies to isolate the effect of invariantization, equivariance, and the number of octic blocks. In particular, we achieved an approximate 40% reduction in FLOPs for ViT-H without sacrificing accuracy, positioning octic ViTs as a strong addition to the catalog of vision architectures.
@@ 31 SQRT2_OVER_4 * (bmd + emh)
A.2 Time complexity
[…]
Consistent with previous results, we find that the equivariant linear layers pay off more with increasing embedding dimension. The extra computations needed for the non-linearity give only a few percent overhead while the quadratic savings from the linear layers provide substantial performance gains.
[…]
Model $C=256$ $C=512$ $C=1024$ $C=2048$
GELU 44.8 82.8 155.8 302.6
[…]
MLP 128.4 324.2 1053.3 3644.9
[…]
Appendix B Arithmetic Intensity
The arithmetic intensity measures FLOPs per transferred byte and can be compared with the FLOPs per bandwidth of a given device to obtain a bound on the maximum achievable throughput (Austin et al., 2025). The arithmetic intensities are $\frac{2BCF}{P(BC+CF+BF)}$ and $\frac{2BCF/5.33}{P(BC+CF/8+BF)}$ for the standard and octic linear layers, respectively, where $B$ is the batch size in tokens, $C$ is the input dimension, $F$ is the output dimension and $P$ is the precision in bytes. This means that the octic and ordinary layers scale differently. At large scale, not only FLOPs are improved by octic layers but also arithmetic intensity. For instance, for $B=196$ (one image worth of tokens), $P=2$ and $F=4C$ (a typical MLP expansion factor) one can calculate that ordinary linear layers have higher arithmetic intensity up to $C\approx 3200$, whereas octic linear layers have higher arithmetic intensity at larger dimensions. For the experiments in this paper, we are not able to scale to such large dimensions, but still get throughput benefits due to savings in FLOPs, as shown in Table 1.
Appendix C Visualizations
We visualize the action of the octic group on images and on $\rho_{\text{iso}}$-features in Figure 5.
(a) PatchEmbed filters.
Figure 5: (a) PatchEmbed filters from a trained network. More filters are shown in Figure 6.
(b) Octic action on patchified images.
(c) Octic action on $\rho_{\text{iso}}$-features.
C.1 Learned Filters
The PatchEmbed layer contains filters mapping the input from 3 channel dimensions to $C$ embedding dimensions. To illustrate the learned filters, we take inspiration from Dosovitskiy et al. (2021) and visualize the first 16 principal components. In contrast to regular ViTs, we have six different learned filters corresponding to the five irreps. Four for the one-dimensional irrep and two for the E irrep (due to its multiplicity). The results are illustrated in Figure 6. Interestingly, the learned filters look qualitatively different between the two learning methods. This is similar to how the learned filters of baseline DINOv2 and DeiT III look qualitatively different. DINOv2 training appears to produce more high frequency patterns while DeiT III gives clearer patterns. For the invariant irrep (A1), it appears that the DeiT III training produces a spherical pattern.
@@ C.1 Learned Filters
Appendix D Invariantization
There are multiple options to produce $\mathrm{D}_{8}$ invariant features, i.e. mapping tokens of type $\rho_{\text{chan}}=\frac{C}{8}\rho_{\text{iso}}$ to features of type $C\rho_{\text{A1}}$ (here denoted in short as invariantization). We let $\psi$ be a function mapping from features of type $\frac{C}{8}\rho_{\text{iso}}$ to features of type $\frac{KC}{8}\rho_{\text{A1}}$ for some $K$ which can be larger or smaller than $8$. These $KC/8$ dimensions are then mapped through a small MLP to $C$ dimensions again.
@@ Appendix D Invariantization
The triple correlation method (Sanborn & Miolane, 2023; Kakarala, 2012) extracts a complete set of third order homogeneous polynomial invariants from a signal over $\mathrm{D}_{8}$. We computed a basis for all third order invariant homogeneous polynomials using Macaulay2 (Grayson & Stillman, ; Ferraro et al., 2024) and use the basis elements as invariant. Here, $K=15$.
@@
We follow Cahill et al. (2024) and implement a version of their max filtering invariant. For this, we have a set of $2C$ learnable $C$-dimensional tokens $\mathbf{y}\in\mathbb{R}^{2C\times C}$, and the $2C$ invariants are given by
$\psi(x)=\oplus_{k=1}^{2C}\max_{g\in\mathrm{D}_{8}}\langle\mathbf{y}_{k},\rho_{\text{chan}}(g)x\rangle.$ (16)
@@
Similar to the max filtering approach, we implement a canonisation where we have a single learnable $C$-dimensional reference token $y$ and compute the $C$-dimensional invariant as
@@
We conduct a study of the effect of these different invariantization methods. A priori, max filtering and canonisation should be more expressive than the others as they are the only invariants considered here that are able to preserve the relative phase information coming from different phase in different copies of $\rho_{\text{iso}}$. We train $\mathrm{D}_{8}(\text{ViT-L/16})$ on ImageNet-1K following the DeiT III recipe. The results are presented in Table 5. The conclusion is that the simple power spectrum invariant works well, and so we select it as our invariantization of choice in the remainder of the experiments.
@@
Appendix E Experimental Setting
E.1 DeiT III
We train for 400 epochs on ImageNet-1K with an effective batch size of 2048 following Touvron et al. (2022) and Bökman et al. (2025). We compare to the figures reported in the respective papers and thus only train the octic ViTs. The training recipe includes heavy data augmentation (e.g. cutmix, mixup and color jitter) and uses the deprecated NVIDIA Apex library. Training is done in mixed precision with the lamb optimizer. The most important hyperparameters are summarized in Table 6(a). For more details, we refer to (Touvron et al., 2022; Bökman et al., 2025) and the official repo used for reproduction https://github.com/facebookresearch/deit.
@@ E.1 DeiT III
Init layer scale $10^{-5} 1.1$$10^{-5}$
@@
E.2 DINOv2
We closely follow the implementation in the original paper, only modifying to train in BF16 instead of FP16 for greater stability, and follow the same evaluation protocol for classification. Lacking an official reproduction of the segmentation protocol in DINOv2, we opt for the evaluation protocol created by Darcet et al. (2025) for semantic segmentation on ADE20K and VOC2012. In contrast to the original DINOv2 paper, we decide to limit our study to ViT-L and ViT-H, the latter of which was not included in the original paper. We opt for the larger patch size of $P=16$ for all our DINOv2 models to save computational resources. Note, this is the reason why we report fewer FLOPs for our DINOv2 ViT-H models than their DeiT III counterpart (which use $P=14$).
@@ E.2 DINOv2
We extend the evaluation of DINOv2 (trained on ImageNet-1K) to two popular classification datasets and report the performance in Table 7. We find, once again, that our networks achieve similar or better performance than the baseline while using substantially fewer FLOPs.
Table 7: DINOv2 additional evaluation. We further evaluate the frozen DINOv2 features by classification accuracy on iNaturalist2021 (Van Horn et al., 2021) and Places365 (Zhou et al., 2017a).
Model FLOPs iNaturalist2021 $\uparrow$ Places365 $\uparrow$
($\times 10^{9}$) linear $k$-NN linear $k$-NN
ViT-H/16 127.7 81.7 81.0 34.7 30.6
$\mathcal{I}_{8}(\text{ViT-H/16})$ 77.7 81.9 80.9 33.9 29.2
$\mathcal{H}_{8}(\text{ViT-H/16})$ 77.5 82.2 81.4 35.1 31.1
ViT-L/16 61.9 80.9 80.5 33.2 28.4
$\mathcal{I}_{8}(\text{ViT-L/16})$ 38.5 81.2 80.4 32.6 28.0
$\mathcal{H}_{8}(\text{ViT-L/16})$ 37.7 81.3 80.8 33.6 29.4
E.3 DinoBloom
We investigate the performance of our invariant model on white blood cell classification. We follow the procedure of DinoBloom (Koch et al., 2024) and report our results for ViT-L in Table 9. We find that the invariant model $\mathcal{I}_{8}(\text{ViT-L/16})$ outperforms the baseline on most evaluated metrics. We tried evaluating on a rotated test set and found negligible change in performance for the baseline (the invariant model inherently has no change in performance, similar to the last column of Table 2).
@@ E.3 DinoBloom
$\mathcal{I}_{8}(\text{ViT-L/16})$ 78.1 78.0 61.3 83.5 83.9 55.7 85.0 85.2 61.9
E.4 General Settings
@@
We utilize the PyTorch (Paszke et al., 2019) and the timm (Wightman, 2019) libraries for our experiments. We run the same versioning as our benchmarks. For all other experiments, we use Python 3.11.9 and PyTorch 2.6.0 with CUDA 11.8.
@@
The model sizes referred to in the paper adhere to the standard terminology used by Wightman (2019); Dosovitskiy et al. (2021). If we denote the shape by a tuple of (depth, width, attention heads), ViT-L has shape (24, 1024, 16) and ViT-H has shape (32, 1280, 16). Both use MLP dimension four times the size of the embedding dimension (commonly referred to as MLP ratio).
@@
Throughput and peak memory are measured on a single A100-80GB GPU with batch size fixed to 64 using torch.compile, FlashAttention (Dao, 2024), and mixed precision. The throughput only measures forward passes with no gradients. Moreover, we utilize 10 warm-up iterations and then average over 100 runs (Bökman et al., 2025). Peak memory is measured with PyTorch’s device memory allocation monitor.
@@
We count the number of FLOPs using fvcore.nn.FlopCountAnalysis https://github.com/facebookresearch/fvcore. FLOPs are normalized with respect to the batch size (i.e. we measure FLOPs/image). We acknowledge that the term FLOPs often leads to confusion. We adopt the terminology of prior work (Touvron et al., 2022; Bökman et al., 2025) and the fvcore library for FLOPs, though, strictly speaking, this refers to MACs (as a factor of two is omitted).
E.5 Compute Resources
Table 10 provides information on the computing resources required for the main experiments. The largest single experiment (i.e. not accounting ablation studies) used $32\timesA$100-40GB for 58 hours. Failed or discarded results are not included. In total, we used around 20k A100-40GB hours for our main results.
@@ E.5 Compute Resources
DeiT III $\mathcal{I}_{8}$(ViT-H/14) $32\timesA$100-40GB 58h 1856
DeiT III $\mathrm{D}_{8}$(ViT-H/14) $32\timesA$100-40GB 57h 1824
DeiT III $\mathcal{H}_{8}$(ViT-H/14) $32\timesA$100-40GB 55h 1760
DeiT III $\mathcal{H}_{8}$(ViT-L/16) $16\timesA$100-40GB 39h 624
DeiT III $\mathcal{I}_{8}$(ViT-L/16) $16\timesA$100-40GB 37h 592
DeiT III $\mathrm{D}_{8}$(ViT-L/16) $16\timesA$100-40GB 36h 576
Ablation: Invarisation $\mathcal{H}_{8}$(ViT-L/16) $16\timesA$100-40GB 240h 3840
Ablation: Hybridisation $\mathcal{I}_{8}$(ViT-L/16) $16\timesA$100-40GB 122h 1952
Ablation: Hybridisation $\mathcal{H}_{8}$(ViT-L/16) $16\timesA$100-40GB 120h 1920
DINOv2 $\mathcal{H}_{8}$(ViT-H/16) $32\timesA$100-40GB 33h 1056
DINOv2 $\mathcal{I}_{8}$(ViT-H/16) $32\timesA$100-40GB 30h 960
DINOv2 $\mathrm{D}_{8}$ (ViT-L/16) $32\timesA$100-40GB 23h 736
DINOv2 ViT-H/16 $32\timesA$100-40GB 21h 672
DINOv2 $\mathcal{I}_{8}$(ViT-L/16) $32\timesA$100-40GB 19h 608
DINOv2 $\mathcal{H}_{8}$(ViT-L/16) $32\timesA$100-40GB 19h 608
DINOv2 ViT-L/16 $16\timesA$100-40GB 24h 384
DinoBloom $\mathcal{I}_{8}$(ViT-L/16) $8\timesA$100-40GB 46min 6
DinoBloom ViT-L/16 $8\timesA$100-40GB 37min 5
@@ Appendix F Licenses
All our models are trained on ImageNet-1K (Deng et al., 2009) which has a custom non-commercial license (see https://www.image-net.org/download.php). We also evaluate ADE20K, VOC2012, and BMC. For the licenses, see https://ade20k.csail.mit.edu/terms/index.html, http://host.robots.ox.ac.uk/pascal/VOC/voc2012/, and https://www.cancerimagingarchive.net/data-usage-policies-and-restrictions/, respectively.
@@
The efficient octic layers are our original work and will be licensed under Apache License 2.0 following the code release. The training pipelines are from DeiT III (Touvron et al., 2022) and DINOv2 (Oquab et al., 2024), which are also under Apache License 2.0. For details about the Apache License 2.0 we refer to https://www.apache.org/licenses/LICENSE-2.0.